OpenStudy (anonymous):

Choose the correct graph of the following system of inequalities: 3x + 5y < 10 x – 4y≤ 12

OpenStudy (zepp):

Isolate y first.

OpenStudy (zepp):

3x + 5y < 10 5y < 10-3x y < (10-3)/5

OpenStudy (anonymous):

this site really helps... https://www.desmos.com/calculator

OpenStudy (zepp):

x – 4y≤ 12 4y ≤ x-12 y ≤ (x-12)/4

OpenStudy (anonymous):

idk.how.to.do.it.on.the.graph.afta.i.isolate.the.y

OpenStudy (zepp):

@uj1024 It really doesn't..... when she's taking a test :(

OpenStudy (zepp):

Okay, we have \(\large y < \frac{(10-3)}{5}\) and \(\large y ≤ \frac{(x-12)}{4}\)

OpenStudy (zepp):

The '<' sign indicate that the inequality doesn't include the line.

OpenStudy (anonymous):

where.the.x.go.in.the.first.one?

OpenStudy (zepp):

If we drag the slope down. We'll get: \(\large \frac{-3x+10}{5} \rightarrow -\frac{3}{5}x+2\) Sorry, the x completely disappeared D:

OpenStudy (zepp):

and \(\large \frac{x-12}{4} \rightarrow \frac{1}{4}x-3\)

OpenStudy (anonymous):

ook.i.got.that.part.i.think

OpenStudy (zepp):

So we have \(\huge y<-\frac{3}{5}x+2\) and \(\huge y≤\frac{1}{4}x-3\)

OpenStudy (zepp):

The first inequation's slope is negative, therefore it's going down, and the other one is going up.

OpenStudy (anonymous):

its.the.last.one.or.the.second.one.but.i.dont.understand.how.to.graph.it

OpenStudy (zepp):

Now, if y is smaller than the function itself, it mean that the shadowed zone should be all under both inequations.

OpenStudy (zepp):

 and the function should be dotted, since it's a > sign.